OpenStudy (johnnydicamillo):

Find the most general AntiDerivative of f(x) = x -3

OpenStudy (anonymous):

use the power rule backwards

OpenStudy (anonymous):

for example, the anti derivative of \(x\) is \(\frac{x^2}{2}\)

OpenStudy (johnnydicamillo):

right but do I need to write is as 1/2x?

OpenStudy (anonymous):

in general the anti derivative of \(x^n\) is \(\frac{x^{n+1}}{n+1}\)

OpenStudy (johnnydicamillo):

\[\frac{ 1x^2 }{ 2 }\]? or will just x^2/2 be okay

OpenStudy (anonymous):

jeez not to be annoying, but when you get to calc, it is time to drop the "invisible 1" right?

OpenStudy (johnnydicamillo):

well the answer in the book put is there, so I did not know if it was significant.

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

wow i guess it was trying to "show all steps" as they say yes, just go with \[\frac{x^2}{2}-3x+C\]

OpenStudy (anonymous):

don't forget that \(+C\) since it asked for "most general" one

OpenStudy (johnnydicamillo):

right, thanks!

OpenStudy (anonymous):

yw